OpenStudy (anonymous):

What is the combined area of the two bases of the following cylinder? 452.16 cm2 904.32 cm2 113.04 cm2 226.08 cm2

OpenStudy (anonymous):



OpenStudy (nirmalnema):

what u need area of whole cylinder or only of bases

OpenStudy (anonymous):

Okay, how do you find that?

OpenStudy (nirmalnema):

area of base = area of circle pi *r^2 but there are 2 circles so 2(pi*r^2) r=d/2 d=12 so r= 6 (2*3.14*6*6) =226.08

OpenStudy (anonymous):

Okay, so now what?

OpenStudy (nirmalnema):

what?? 226.08 is the combined area of two circles. what else u need

OpenStudy (anonymous):

Sorry brain fart!

OpenStudy (nirmalnema):

lol.ok no problem

OpenStudy (anonymous):

What about the surface area of that same figure?

OpenStudy (nirmalnema):

total surface area or curved surface area??

OpenStudy (anonymous):

Total.

OpenStudy (nirmalnema):

total surface area = curved surface area + area of two bases csa=2pi *r*h =2*3.14*6*15=565.2 565.2+area of two bases =565.2+226.08 =791.28

OpenStudy (anonymous):

You're so much help! Thank you!

OpenStudy (nirmalnema):

no problem.. your welcome

OpenStudy (anonymous):

Do you know anything about deminsions? Im having trouble understanding them.

OpenStudy (nirmalnema):

yaa sure i will try

OpenStudy (anonymous):

If the dimensions of the following cylinder are tripled, what factor will the volume change by? 3 9 27 81 Its the same figure as above

OpenStudy (anonymous):

Dont we need to find the volume too?

OpenStudy (nirmalnema):

no. volume of cylinder = pi*r^2*h if r and h is tripled then new formula is pi*(3r)^2*(3h) so the volume is increased = pi*9r^2*3h =27 times

OpenStudy (anonymous):

So to find the volume you would just triple everything right? Or no?

OpenStudy (anonymous):

@driftracer305 If the dimensions of the following cylinder are tripled, what will be the volume of the new cylinder?(same figure)nI dont understand it at all!

OpenStudy (anonymous):

Wait, got it!

OpenStudy (driftracer305):

wat u got?

OpenStudy (driftracer305):

it shld be 27 times....... he was right...... all good